OpenStudy (anonymous):

Also one more Maria and paul collect marbles. Paul has one third as many marbles as maria. Together they have 48 marbles . How many marbles does Maria have?

OpenStudy (luffingsails):

Let X = number of marbles that Maria has. Then (1/3)*X = number of marbles that Paul has. "Together they have 48 marbles" means that: X + (1/3)X = 48 Simplify and solve for X. Does that help?

OpenStudy (anonymous):

So ?

OpenStudy (luffingsails):

So, can you take it from there?

OpenStudy (anonymous):

Can you help me a little more or give me the answes haha il give good answer on you ?

OpenStudy (luffingsails):

I'm not interested in a medal. lol... I am interested in you getting how to solve the problem. What is the result of: \[\frac{3}{3}x + \frac{1}{3}x\]

OpenStudy (anonymous):

10?

OpenStudy (luffingsails):

Just add the coefficients. \[\frac{3}{3}x+\frac{1}{3}x = x(\frac{3}{3}+\frac{1}{3}) = ?\]

OpenStudy (anonymous):

what do you mean?

OpenStudy (luffingsails):

I'm asking you to simplify the equation. You do that by adding the coefficients: 1 + 1/3 = (3/3) + (1/3)

OpenStudy (anonymous):

4/6